You have 80 grams of a radioactive kind of tellurium. How much will be left after 8 months if its half-life is 2 months?





1
Expert's answer
2022-03-30T11:16:01-0400

"N(t)=N_0(\frac{1}{2})^{\frac{t}{t_{1/2}}}=80\cdot(\frac{1}{2})^{\frac{8}{2}}=5\space g."